Calculate the total moles of atoms of each element present in 122.5 g of KClO3.
1
Expert's answer
2018-08-03T05:09:42-0400
M(KClO3)= 39+35,5+3*16) = 122.5 g/mol
ν(KClO3) = 122.5 g / 122.5 g/mol = 1 mol
Consequently, 122.5 g KClO3 content 1 mole of of K and Cl and 3 moles of O.